SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

BaBar homepage  CALORIMETER CALIBRATION MEETINGS

back to: EMC Calibration page

The calorimeter calibration meetings take place on WEDNESDAYS (beginning with Oct 6, joined with a EMC operations meeting), weekly at 9:30 a.m. PST (5:30 p.m. UK, 6:30 p.m. Central Europe) in the Group A conferennce room (CLA, B188). If you wish to be connected by phone call the DCS conferencing system at +1-510-647-3480. If you have a touch-tone telephone, you may then connect directly to the teleconference by pressing 1 and dialing the passcode 104530 followed by the "number sign" ("#") key. Otherwise, you may request the "EMC software meeting moderated by Helmut Marsiske". Occasionally, when we run out of DCS reservations, we may use the STS conference center at +1-973-321-2000, passcode 107969#. 

A postscript copy of your transparencies should be made prior to the meeting and posted to this page. To do this do the following: 

1. Name your .ps file as XXXDDMMMYYII.ps where XXX=gen,rad,norad,src,pls,ele,pi0,mip (i.e the type of calibration), DDMMMYY is date of meeting and II are the initials of the author eg. nonrad25mar99rm.ps. 

2. Copy the file to the following area: 

$BFROOT/www/Detector/Calorimeter/software/calib_mtg 

3 Send mail to jessop@slac.stanford.edu with the name of the postcript file in the title 

Standing agenda items (and links to working web pages) are:

  • General Calibration Issues 
  • Non-Radiative Bhabha calibration 
  • MIP Calibration 
  • Source calibration
  • Light Pulser Calibration 
  • Radiative Bhabha calibration
  • Electronics Calibration 

  • 25.March 99 1.April 99 8.April 99 15.April 99 22.April 99
    29.April 99 6.May 99 13.May 99 20.May 99 27.May 99
    3.June 99 10.June 99 17.June 99 1.July 99 22.July 99
    29.July 99 5.August 99 19.August 99 2.September 99 9.September 99
    16.September 99 23.September 99 30.September 99 6.October 99 13.October 99
    27.October 99 3.November 99 10.November 99 17.November 99 24.November 99
    1.December 99 15.December 99 5.January 00 12.January 00 19.January 00
    26.January 00 2.February 00 9.February 00 23.February 00 1.March 00
    8.March 00 15.March 00 22.March 00 29.March 00 12.April 00
    19.April 00 26.April 00 3.May 00 12.May 00 17.May 00
    24.May 00 31.May 00 7.June 00 21.June 00 2.August 00
    9.August 00 16.August 00 23.August 00 6.September 00 13.September 00
    20.September 00 4.October 00 18.October 00 8.November 00 15.November 00
    29.November 00 6.December 00 20.December 00



    25. March 99 :



    1. April 99 :

    • QC light yields Light yields measured by scanner in QC process have been implemented in barrel. Effect of implementing QC crystal light yields is to improve calibration to 7%. Agrees with Hiro's work. Conclusion is that QC light yields can be used as first calibration to 7%. postscript Also that constants in database are implemented correctly (this is first calibration we have had). - Leif 

    • End cap QC light yields will be available in a few days.Important because no MIP for endcap and QC gives 7% initial calibration - Roger. 
    • Non-rad bhabars: To extract single crystal constants from clusters need MC simulation. Previously done with DIRC in place (400K events) now redoing without DIRC. Also with variable binning of rec impact point to overcome clustering of statistics at crystal centroid. - Ralph 
    • MIPs with tracks. Swim tracks through crystals and use de/dx to calibrate. Preselect events with good clusters so don't waste CPU on swimmer with bad clusters. Problems with digi timing cut now solved. Tuning up other cuts. Production should begin in a few days. - Francesca 
    • MIPS without tracks postcript Use back to back single crystal clusters to find sample of muons where muon has deposited most of energy in one crystal. Then use to calibrate. Can calibrate approx 2000 crystals to better than QC light yield (2-5%). Advantage is that it is quick and requires only a few days of processing. Production of these constants will start this week. - Leif 
    • Pi0 Photon selection cuts are being tuned by Elias - results in two weeks. - Roger 

    • Learning how to use spatial and temporal databasesa - Akram 
    • Electronics Will attempt full electronics calibration in next few days. 
    • Source Generator is scheduled to arrive on April 15th. It is probably too late for us to be able to start with new generator hence we will start with old generator (factor of 7 less rate than designed) and try to debug and do as well as we can with this. System is being reassembled and should be able to start running Monday. - Matt 
    • Source DAQ . Problems with dataflow require expert assistance. Meeting with experts today. - Frank 
    • Light Pulser I/O board for calibration has been fixed at SLAC - a problem with a NIM output. Roland will come to SLAC when he finds his passport to commision the calibration system. Phil is back in UK for two weeks. Francesca will take over light pulser responsibilities from Phil when she arrives May 1 - Steve Playfer. 

    • No report from Germany. 
    • Rad Bhabhas Still Problems with fit converging when use lagrange multipliers. Meet tomorrow with expert (Bill Dunwoodie) to see if we can solve problems - Johannes 



    8. April 99 :

    • General Issues . MIP and LY constants should be in place by May 1. New cosmics runs April 20th - April 28th without magnet. April 29th May 3rd with magnet. EmcOepDiag changes location from EmcOepMon to tc_to_digi package. 
    • QC light yields Roger has still not supplied the End cap light yields - is it secret information ? 
    • Non-rad bhabars: Not enough statistics to study effect of DIRC on extraction of single crystal constants.Perhaps next week. - Ralph 
    • MIPs with tracks. Production underway should take one week - quicker than expected ! - Francesca 
    • MIPS without tracks First results of production of constants - Leif 
    • Pi0 No report. Akram away learning about objectivity. 
    • Electronics Plots Both hardware and software are now all working so the first attempts at a full calibration have taken place. Six constants are generated four gains for the the different ranges, a common offset and a ratio corresponding to the different capacitances used for charge injection into the different ranges. There is a non-linearity in the low range of the x1 range due to crosstalk from the different channels. This is a hardware bug that was fixed but not in time for the preamp production run. Should not be a problem however because the auto-ranging means that we never use this anyway. Constants had been generated with auto-ranging off. Ultimately will be generated wit it on. More work is necessary to refine the proceedure. -Jordan 
    • Source DAQ on test stand now working again. Plumbing should be installed by next Thursday - so will begin 3 day attempt at getting a full calibration run. New neutron source may arrive 15th and every effort will be made to get it in place. - Frank/Matt 
    • Light Pulser 
    • Rad Bhabhas Fit still not converging. - Johannes 



    15. April 99 :

    • General Issues
    • QC light yields No response to repeated requests to Manchester to get light yields for endcap. Conclude that we will not get these and therefore have no starting calibration point for endcap. 
    • Non-rad bhabars: Ralph has arrived at SLAC and is working on implementing calibration in the new OEP release. Almost complete.
    • MIPs with tracks. Production is complete. James and Francesca have to organise how to do the fits to generate constants. First set of constants next week - Francesca 
    • MIPS without tracks No Report. Leif at conference in Berlin. - Leif 
    • Pi0 No progress.. 
    • Electronics Hardware bug that caused non-linearity small non-linearity in overlap regions is not a problem as these regions are not used when auto-ranging. Possibly a full calibration by Monday - assuming dataflow system is working. 
    • Source Software works and hardware is back together in IR2. Upgrades in dataflow appear to have made things faster. Attempts at first run continue tonight and in near future. - Frank/Matt 
    • Light Pulser Plots Board fixed but awaits connector. Phil will return it to SLAC next week. He and Martin will begin commissioning system on Wednesday.
    • Rad Bhabhas Plots Much progress. The fit converges - rapidly!. - Johannes 



    22. April 99 :

    • General Issues
    • End Cap QC Light Yields: Light yield numbers have arrived and will be put into the database by Michael George with help from Peter Kim
    • MIPs with tracks. First constants.- Francesca/Jim 
    • MIPS without tracks Comparing MIP with tracks and MIP without tracks leif finds agreement in backward barrel but not the forward barrel where it is off by 10%. 
    • Pi0 Code is implemented in OPR but not yet fully tested.Akram 
    • Electronics No report - Jordan/Peter 
    • Source Software is all working. Live time has increased from 2% to 6% making the calibration 3 times faster. Max possible in DAQ is 10%. Old generator has problems with HV supply and we are working with vendor to try to fix them. Safety clearance to run system has expired so wont have next chance to run until next tuesday.Frank/Matt 
    • Light Pulser Calibration board and connectors have arrived with Phil. Waiting for code to program Xylinx chips from Roland. Expect in the next two days. Should now have all parts of system in working order except the capacity to chance light pulser intensity. I.e all components necessary to monitor light yield chances due to radiation are inplace and should be working soon. Martin Dickopp will review software status next week. Phil/Martin/Roland
    • Rad Bhabhas Have started to update code but have not had chance to test and see if chisquared problem has gone away in new release. Johannes 
    • Non-rad bhabars: No report



    29. April 99 :

    • MIPs with tracks. plots Constants from Febuary Cosmic Run. First constants show an improvement over previous estimated constants as evidenced by a 25% narrowing of MIP peak - long tails. Automated fit proceedure has no quantification of errors associated with fit so bad fit may cause tails.- Francesca/Jim 
    • MIPs with tracks. Status of April cosmic run constants analysis. First runs now being analyzed with new swimmer. 2 million events with track triggers- Hiro 
    • End Cap QC Light Yields: Speadsheets from Roger have problems. Will be fixed and then rescaled to give a calibration number. - Michael George
    • MIPS without tracks Febuary constants still in progress- some technical problems with running MC to understand the effect of staggering. First look at April cosmics - very inefficient EMT trigger for usuable back to back cosmic clusters (2%). L3 ups to 20%.
    • Pi0 OPR implementation. Akram 
    • Electronics Status Jordan/Peter 
    • Source Old generator gutted. Awaiting new generator from Russia. New generator can be installed in an 8 hour shift when it arrives. - Matt
    • Light Pulser - hardware Calibration board now working. Waiting to do full system test Martin K./Roland
    • Light Pulser - software Software status report Software all working and making correct plots. Some problems with missed events. Martin D.
    • Rad Bhabhas plots Moved to 7.12.8. Cluster and track errors look incorrect. Believe fixed in 8.1.1. Move to this for next week.. - Johannes 
    • Non-rad bhabars: No report - Ralph



    6. May 99 :

    • MIPs with tracks. Febuary cosmics: chi squared of MIP fits looks good. chi squared versus Emeas/Epredicted shows no correlation hence long tails not associated with bad fits. Systematic uncertainty due to fitting with peak and mean needs to be understood. Comparison of LY calibs with MIP is linear but does not intersect at 0. Looking at the ratio shows a gaussian with a 9% width and a high side tail. Expect gaussian due to measurement errors not tail. Attempt to correlate tail with crystal properties show no correlation. If remove crystals in tail and then plot ly versus MIP again then intersects closer to 0 so tail causes error in calib. More work is underway to understand this. Also working to understand systematics due to using the peak instead of mean as calibration point. - Francesca/Jim/Hiro 
    • MIPs with tracks. Status of April cosmic run constants analysis. Preliminary looks at data indicate that the range change problem that causes two peaks in digi distribution is gone. End cap data now available - awaits analysis. - Francesca/Jim/Hiro 
    • MIPs without tracks. Constants from Febuary. MC carlo dataset of MIPS in forward endcap is now complete and should allow generation of a set of constants within a few days. . April data: L3 boost efficency to 20% levels as opposed to 2% so we are now taking useful data.- Leif 
    • MIPs in OPR. Fitting implemented. Needs L3 MIP selector. Jim wil write this and Hiro will take care of data checking at this end. - Jim 
    • End Cap QC Light Yields: Roger will produce final numbers Friday.
    • Pi0: Proceedure is implemenetd and runs on MC events: Akram 
    • Electronics: Constants now for entire detector. Range change non-linearities still exist but effects shold be minor. Differences in preamp gains not changed in electronics (just linearity) - will be taken out by light yield calibs. Jordan/Peter 
    • Source: No word of neutron gun arrival. Continued software development. - Matt/Frank
    • Light Pulser: - Now pulsing endcap. 1 crystal is masked off (by Jordan) and one ROM not yet checked out (40 crystals). 4 noisy crystals in the rest."Energy" plots shown for barrel and endcap . The strange peak occurs because peak of distribution occurs at range change in electronics. Unclear whether this affects ability to monitor rad damage. Work continues to make entire system working - then to establish its stability with a set of calibrations over a period of time. Martin K/Martin D/Phil C/Francesca 
    • Rad Bhabhas: Still track errors in 8.1.1. Fit errors worked on by Matt while Johannes concentrates on getting OPR working by next week. Probably will not be ready for data next week.- Johannes 
    • Non-rad bhabars: Complete executable works on MC simulated events. Executable needs to be broken down to run in multinode OEP enevironment - should happen this week with help from John Bartelt. Should be ready to calibrate with data next week. - Ralph



    13. May 99 :

    • MIPs without tracks. Qualification plots for April cosmics - Leif 
    • MIPs with tracks. systematic errors - Jim Febuary cosmics: Qualification plots for cosmics that went in Monday 10th May. Estimate of accuracy of constants - Francesca/Jim/Hiro 
    • MIPs with tracks. Status of April cosmic run: executable with Elf is now ready (for Jim: we needed to pass to elf because of the DCH, i.e. of tracks) with the things needed for the off--line calibration added to the EmcCalMinI package. I sent few jobs during the night but they aborted (it's not a crash in the program but a failure in Bronco during the night in accessing the database. In fact, when I tested jobs interactively, but of course with only few events, they were ok). Optimistically speaking, we could have new calibration constants for mips w/tracks at the beginning of next week (Monday...) and then check them with Leif's ones. I would like to emphasize that April/May data look much better then January run, so results should be better than before. - Francesca/Jim/Hiro 
    • MIPs in OPR. Status of L3 MIP selector. - Jim/Colin 
    • End Cap QC Light Yields: Plot of endcap QC light yields - Roger/Micheal/Hiro
    • Pi0: I have commited a tag for EmcCalPi0 V00-01-01. Also Roger has started to look at the numbers that he produced at the Feb. Collab. in greater detail. I have been trying to run the Calibration datastream (which includes the EmcCalMinI and EmcCalPi0) within Elf and am making progress but have not completed this. Too many shifts and Elf (Opr) is not in a good shape at yet, too many fixes! : Akram 
    • Electronics: Status - Jordan/Peter 
    • Source: Status. - Matt/Frank
    • Light Pulser: - plots Plots to demonstrate all system is working and calibrated. Plots to show stability of system over time to demonstrate that it can monitor radiation damage. Martin K/Martin D/Phil C/Francesca 
    • Rad Bhabhas: Are we ready for data ? - Johannes 
    • Non-rad bhabars: Are we ready for data ?. - Ralph



    20. May 99 :

    • MIPs with/without tracks. New set of constants now available. Combines three methods 1) Without tracks - April Data 2) With Tracks Febuary data 3) QC constants. This is necessary because 1) only in central barrel due to back to back requirement 2) limited by cosmic distribution. Constants go in preference 1,2,3 i.e where constants of type 1) are available they are used. Constants of type 2 are corrected for mean/peak difference explained 13th May meeting. QC numbers are scaled using the MIP/QC ratio for April data for a set of high statistics good fit crystals (230 crystals). . - Leif/francesca/jim/hiro 
    • Source: Generator has apparently been shipped and is on its way. More work on software in the meantime.. - Matt/Frank
    • Light Pulser: - plots Martin K/Martin D/Phil C/Francesca 
    • Rad Bhabhas: - Johannes 
    • Non-rad bhabars: Still OEP problems related to database access. Ralph returens to Europe but Martin D will be responsible at SLAC - Ralph



    27. May 99 :

    • MIPs without tracks. Systematics studies Generating MIPs with different IP distributions that correspond to the uncertainty in matching MC distributions to data distributions leads to a <3 % change in the MIP. Hence the systematic uncertainty in the calibration is < 3% - Leif 
    • MIPs with tracks. April data The new swimmer is now shown to be working by comparing distributions from the old. Following a check that the emeasured/epredicted is sensible the production of april cosmics will begin Francesca/Jim/Hiro 
    • Pi0:Detailed status report - All elements of the package for calibration are working independently but the enitire package is not quite ready. Expect to be ready to analyse data in 2-3 weeks. Roger 
    • Source: Still no sign of generator.. - Matt/Frank
    • Light Pulser: plots Martin K/Martin D/Phil C/Francesca 
    • Non-rad bhabars: status - Ralph



    3. June 99 :

    • Non-rad bhabas: status - Ralph
    • MIPs with tracks. Calibration from April data is still in progress. Francesca/Jim/Hiro 
    • Light Pulser: plots Martin K/Martin D/Bernd 



    10. June 99 :

    17.June 99 :

    • Non-rad bhabas: status - Ralph
    • MIPs with tracks. Calibration from April data Francesca/Jim/Hiro 
    • Pi0:status-. Akram 
    • Source: - Matt/Frank
    • Light Pulser: plots Martin K/Martin D/Bernd 
    • Radiative bhabha's : - Johannes

    17.June 99 :

    • Non-rad bhabas: status - Ralph
    • MIPs with tracks. Calibration from April data Francesca/Jim/Hiro 
    • Pi0:status-. Akram 
    • Source: - Matt/Frank
    • Light Pulser: plots Martin K/Martin D/Bernd 
    • Radiative bhabha's : - Johannes

    1.July 99 :

    • Final Cosmic Calibration: plots - Francesca 
    • Swimmer Test plots - Jim 
    • Radiative bhabha's:plots - Johannes
    • Non-rad bhabhas: - Leif
    • Non-rad bhabhas: - Tatania
    • Non-rad bhabhas: plots - Jane
    • Non-rad bhabhas: - Ralph
    • Final Cosmic Calibration Francesca/Jim/Hiro 
    • Pi0 status: - Akram 
    • Source: - Matt/Frank
    • Light Pulser: plots - Martin K/Martin D/Bernd 


    22.July 99 :

    • General Status - Colin 
    • Non-Radiative Bhabha - Ralph - first results 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 
    • Radiative Bhabha Calibration status plots- Johannes 
    • Source - First Calibration - Matt pdf/ps
    • Source - First Calibration - Frank html
    • Light Pulser - Martin K 
    • Pi0 status - Akram 


    29.July 99 :

    • General Status - Colin 
    • Constants comparison - Leif ps
    • MC studies - Jane ps
    • Light Pulser - linearity study - Martin K 
    • Electronics linearity check - Peter/Jordan 
    • Non-Radiative Bhabha - The l3 track match problem results 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane ps
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 
    • Radiative Bhabha Calibration status - Johannes 
    • Source - Matt/Frank 
    • Pi0 status - Akram 


    5.August 99 :

    • General Status - Colin 
    • Light Pulser/Electronics linearity - Martin/Peter 
    • Non-Radiative Bhabha: new constants - Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 
    • Radiative Bhabha Calibration status - Johannes 
    • Source - Matt/Frank 
    • Pi0 status - Akram 


    19.August 99 :

    • General Status - Colin 
    • Light Pulser/Electronics linearity - Martin/Peter 
    • Non-Radiative Bhabha: new constants - Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 
    • Radiative Bhabha Calibration status - Johannes 
    • Source - Matt/Frank 
    • Pi0 status - Akram/Sven 


    1.September 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro ps
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Roger/Akram ps
    • Radiative Bhabha Calibration status - Johannes ps
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha: new constants - Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 


    9.September 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro ps
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Roger/Akram ps
    • Radiative Bhabha Calibration status - Johannes ps
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha: new constants - Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 


    16.September 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro 
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Roger/Akram 
    • Radiative Bhabha Calibration status - Johannes 
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha: new constants - Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 


    23.September 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro 
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Roger/Akram 
    • Radiative Bhabha Calibration status - Johannes ps
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha - New constants Ralph/Leif 
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 


    30.September 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro 
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Roger/Akram 
    • Radiative Bhabha Calibration status - Johannes ps 
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha - New constants Ralph/Leif - latest results
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben plots
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 


    6.October 99 :

    • General Status - Colin 
    • Electronics linearity - Peter/Jordan/Hiro 
    • Photon cluster corrections - Sven/Jane 
    • Photon cluster corrections - Klaus/Enrico ps 
    • Photon cluster corrections - Roger/Akram 
    • Radiative Bhabha Calibration status - Johannes ps 
    • Source - Matt/Frank ps
    • Non-Radiative Bhabha - Ralph - change of the calibration constants
    • Non-Radiative Bhabha simple calibration - Urs 
    • Non-Radiative Bhabha simple calibration - Hiro/Dave/Ben 
    • MC studies for bhabha calibration - Jane 
    • Non-Radiative Bhabha simple calibration - Francesca/Steve 




    13.October 99 :





    27.October 99 :

    • Implementation of constants - EmcCalib/EmcDigiCalib - Jim/Sven/Helmut
    • Photon cluster corrections - Klaus/Enrico ps
    • Photon cluster corrections - New plots Roger(for Jim)/Akram 
    • Photon cluster corrections - Sven 
    • Non-Radiative Bhabha - E_meas/E_exp plots - Ralph/Leif 
    • Non-Radiative Bhabha - Francesca/Steve 
    • Non-Radiative Bhabha - Hiro/Dave Summary plot1 plot2
    • Source - EMC Noise Frank / Matt  
    • Electronics calibrations - Jordan 
    • Electronics calibration test with Pulser - Bernd 
    • Noisy and dead channels - Urs 
    • Simulation status, pi0 plot - Christos 





    3.November 99 :

    • Implementation of constants - EmcCalib/EmcDigiCalib - Jim/Sven/Helmut
    • Photon cluster corrections - Klaus/Enrico ps
    • Photon cluster corrections - New plots Roger(for Jim)/Akram 
    • Photon cluster corrections - Sven 
    • Nonuniformity studies - Jenshtml 

    • ^M
    • Non-Radiative Bhabha - E_meas/E_exp plots - Ralph/Leif 
    • Non-Radiative Bhabha - Francesca/Steve 
    • Non-Radiative Bhabha - Hiro/Dave Summary plot1 plot2
    • Source - EMC Noise Frank / Matt  
    • Electronics calibrations - Jordan 
    • Electronics calibration test with Pulser - Bernd 
    • Simulation status, pi0 plot - Christos 
    • Rad. Bhabhas - Johannes ps



    10.November 99 :




    17.November 99 :

    • Implementation of constants - EmcCalib/EmcDigiCalib - Jim/Sven/Helmut
    • Photon cluster corrections - Klaus/Enrico 
    • Photon cluster corrections - Sven 
    • Non-Radiative Bhabha - Ralph 

    • ^M
    • Non-Radiative Bhabha Francesca/Steve 
    • Non-Radiative Bhabha - Hiro/Dave 
    • Source - EMC Noise Matt  /Frank 
    • Electronics/Digital Filtering - Jordan/Frank /Matt 
    • Electronics calibration test with Pulser - Bernd 
    • Simulation status, pi0 plot - Christos 
    • Rad. Bhabhas - Johannes ps
    • Electons - Urs 
    • pi0's - Christos/Sven 




    24.November 99 :

    • Implementation of constants - EmcCalib/EmcDigiCalib - Jim/Sven/Helmut
    • Non-Radiative Bhabha - New constants - Ralph 

    • ^M
    • Non-Radiative Bhabha Francesca/Steve 
    • Non-Radiative Bhabha - Hiro/Dave 
    • Photon cluster corrections from pi0 - Sven 
    • Emc Fibre Swaps - Sven 
    • Photon cluster corrections from rad bhabha - Johannes 
    • Source - EMC Noise Matt/Frank
    • Pulser Linearity Test - Bernd 
    • Electronics/Digital Filtering - Jordan/Frank/Matt
    • Simulation status - Christos 
    • Summary of lepton-id AWG- Urs 
    • Summary of neutrals AWG- Christos/Sven 


    1.December 99 :

    • Implementation of constants - EmcCalib/EmcDigiCalib - Jim/Sven/Helmut
    • OPR Plots to check implementation of constants - Steve
    • Photon cluster corrections function from rad bhabha - Johannes 
    • Photon cluster corrections from pi0 - Sven 
    • Non-Radiative Bhabha - New constants - Ralph 

    • ^M
    • Non-Radiative Bhabha Francesca/Steve 
    • LogPos Studies - Hiro/Dave more studies  
    • Source - Matt/Frank 
    • Pulser Linearity Test - Bernd 
    • Electronics/Digital Filtering - Jordan/Frank/Matt
    • Simulation status -pi0 plot - Christos 
    • Summary of lepton-id AWG- Urs 
    • Summary of neutrals AWG- Christos/Sven 


    15.December 99 :



    5.January 00 :

    • Bhabha calibration with tracks, an update on the tracking- Francesca/Steve
    • E/p and various global calibrators PS-file- Urs
    • Source - Matt 


    12.January 00 :



    19.January 00 :



    26.January 00 :



    2.February 00 :



    9.February 00 :



    23.February 00 :



    1.March 00 :



    8.March 00 :



    15.March 00 :



    22.March 00 :



    29.March 00 :



    12.April 00 :



    19.April 00 :



    26.April 00 :



    3.May 00 :



    12.May 00 :



    17.May 00 :



    24.May 00 :

    31.May 00 :

    7.June 00 :

    21.June 00 :

    5.July 00 :

    2.August 00 :

    9.August 00 :

    16.August 00 :

    23.August 00 :

    6.September 00 :

    13.September 00 :

    20.September 00 :

    4.October 00 :

    18.October 00 :

    8.November 00 :

    15.November 00 :

    29.November 00 :

    6.December 00 :

    20.December 00 :


    Colin Jessop